OpenStudy (anonymous):

need help factoring these 4 proBlems 1) s^2-6s-9 2)-x^5+20x^4-100x^3 3)7x^2-63 4) 3b^2-75

OpenStudy (anonymous):

These are all special forms, or reducible to special forms. By special forms, I mean (a+b)^2, (a-b)^2 or a^2-b^2.

OpenStudy (anonymous):

On the third and fourth ones, factor out the common factor first, then they will be in the form of difference of squares.

OpenStudy (anonymous):

for third one is it 7(x^2-9)

OpenStudy (anonymous):

Actually, I don't think (1) factors at all. I think you might have made a sign error on the constant term. If it is +9, then it will be a perfect square.

OpenStudy (anonymous):

thats what i thought... i didnt mess up so it i think its just you cant do it

OpenStudy (anonymous):

The third one will factor further:\[7x^2-63=7(x^2-9)=7(x+3)(x-3)\]

OpenStudy (anonymous):

The fourth one is similar in pattern to the third.

OpenStudy (anonymous):

oh so i need to factor twice... okay i get that

OpenStudy (anonymous):

what about 2nd one though? i dont get it

OpenStudy (anonymous):

It is a perfect square. Recall \[(a+b)^2=a^2+2ab+b^2\]

OpenStudy (anonymous):

\[-x^5+20x^4-100x^3=-x^3(x^2-20x+100)=-x^3(x-10)^2\]

OpenStudy (anonymous):

oh thaat maeks sense

OpenStudy (anonymous):

Factor out the common terms, and look for known patterns before trying to fiddle too much with your regular factoring strategy. Sometimes, it will save you a bunch of trouble.